How many moles of a KOH solution are there in 375mL of 2.45M KOH, and how many grams of KOH do you have


1
Expert's answer
2022-04-28T14:48:03-0400

n(KOH)=CV=0.375*2.45=0.91875 mol, m(KOH)=0.91875*56=51,45 gram